OpenStudy (anonymous):

A triangle has vertices whose coordinates are A (0, 0), B(5, 7), and C(3, 9). Find the coordinates of the midpoints of each side. Answer in terms of a decimal rounded to the nearest tenth or an improper fraction in simplest form. Midpoint of AB = (,) Midpoint of BC = (,) Midpoint of CA = (,)

OpenStudy (anonymous):

To get the midpoint between two points, just figure out the x that is between their x coordinates and the y that is between the y coordinates. Here's a quick example to show you what I mean. Point A (0,2) Point B (1, 0) Their x coordinates are 0 and 1. Exactly between those is 0.5 Their y coordinates are 2 and 0. Exactly between those is 1 So my midpoint is (0.5, 1)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You're welcome =) Think you can do this one now? Feel free to ask me any questions.

OpenStudy (anonymous):

yeah definately !:)

OpenStudy (anonymous):

Midpoint of AB = (5/2, 6) Midpoint of BC = (6,6) Midpoint of CA = (6,5/2) is this right?

OpenStudy (anonymous):

No, but I'm not sure what you're doing =/

OpenStudy (anonymous):

im using this formula (a+b)/2... given to me in the lesson so idk

OpenStudy (anonymous):

That is fine. (a+b)/2 gives you the average, which will be the point exactly between the two, like I told you to find. But somehow you're doing it wrong because (a+b)/2 will not give you the results you got.

OpenStudy (anonymous):

dang:/ ok

OpenStudy (anonymous):

what would i do for Midpoint of AB =??

OpenStudy (anonymous):

Let's start with just points A and B. Their x values are 0 and 5 What x value is exactly between those two? You could do (a+b)/2 if you need to.

OpenStudy (anonymous):

5/2? thenn right?

OpenStudy (anonymous):

Right. Good. 5/2, which is just 2.5. Now look at the y values and do the same thing. Their y values are 0 and 7

OpenStudy (anonymous):

7/2........... omg yes i did something wrongg

OpenStudy (anonymous):

Yes. 7/2, which is 3.5

OpenStudy (anonymous):

So the midpoint for AB is (2.5, 3.5) or you could write it as (5/2, 7/2)

OpenStudy (anonymous):

Let me see your try for AC next.